# Tony Lindeberg

### v2

@@ 6.2 Additional variabilities involving the temporal domain
Concerning the degree of freedom corresponding to temporal scaling transformations, the corresponding degree of freedom in terms of the temporal scale parameter $\sigma_{t}=\sqrt{\tau}$ is also special in the sense that both the non-causal temporal Gaussian kernel and time-causal limit kernel used for temporal smoothing in the idealized model (4.3) for spatio-temporal receptive fields obey cascade properties over temporal scales, implying that the receptive field responses at any coarser temporal scale can be computed by an additional temporal filtering operation being applied to the receptive field responses at any temporal scale. This means that a vision system could, in principle, choose to only implement the earliest layers of temporal receptive fields at the finest temporal scale and nevertheless have the ability to compute the representations at coarser temporal scales, based on additional temporal smoothing applied to the temporal or spatio-temporal receptive field representations at the finest temporal scales. Thus, irrespective of whether the temporal receptive fields are expanded over the temporal scales, it seems plausible that the vision system should have the ability to compute visual operations corresponding to temporal scale covariance.
@@
In Lindeberg (2023b) Sections 3.2.1–3.2.2 and Lindeberg (2025c) Sections 4.2–4.3, sets of suggestions for further neurophysiological experiments are proposed to investigate these hypotheses in more detail, and to characterize the structure of biological receptive fields in the primary visual cortex with respect to the influence of parameters of the receptive fields corresponding to the different degrees of freedom of the 4 main types of geometric image transformations. Complementary suggestions for further research in relation to the influence of geometric image transformations on early vision are also outlined in Lindeberg (2025a) Section 8.1.
@@ 7 Summary and discussion
This theory is based on idealized models of visual receptive fields in terms of combinations of smoothing with spatial smoothing kernels of the form (26) or spatio-temporal smoothing kernels of the form (25) with scale-normalized spatial and temporal derivatives according to Section 5.3. In Sections 5.4 and 5.5, we have described how the resulting idealized models of spatial or spatio-temporal receptive fields obey provable covariance properties under compositions of spatial scaling transformations, spatial affine transformations, Galilean transformations and temporal scaling transformations. Specifically, we have in Section 6 considered the hypothesis about whether the receptive fields of simple cells in the primary visual cortex ought to have their shapes expanded with regard to the degrees of freedom of the basic types of geometric image transformations that occur in the image formation process.
By postulating that the responses of idealized models of receptive fields in terms of scale-normalized spatial and temporal derivative operators are to be possible to match between the image domains before and after the geometric image transformations, we have predicted a set of variabilities over (i) spatial scaling transformations, (ii) image rotations, (iii) the degree of spatial elongation of the receptive fields, (iv) a 4:th spatial degree of freedom, (v) Galilean transformation over image space-time and (vi) temporal scaling transformations. We have considered the plausibility of covariance properties of the either purely spatial or joint spatio-temporal receptive fields with regard to these 7 degrees of freedom (the Galilean transformation comprises 2 degrees of freedom), in view of neurophysiological evidence and structural properties regarding how populations of receptive field responses can be computed based on structural properties of the families of receptive fields under variabilities over the filter parameters. In the absence of sufficient neurophysiological or psychophysical evidence to firmly state whether the predicted properties would hold in the primary visual cortex of higher mammals, we have in Section 6.3 pointed to ideas for future research to characterize these properties in more detail.
@@
  • the spatial scaling factor $S_{t}$ corresponds to the inverse depth $1/Z$ if the affine transformation matrix $A$ in the composed monocular image transformation (5) is normalized in such a way that the affine transformation matrix $A$ reflects a scaled orthographic projection model,
@@
Hence, a vision system that is able to extract these parameters of the geometric image transformations based on processing and comparing populations of receptive field responses, should in principle have the ability to compute direct cues to the structure of environment, directly from established matching relations over the receptive field responses between different views, or in relation to a learned memory of receptive field responses from previous views. Thereby, certain functionalities of the vision system could be formulated directly in terms of the parameters of the primitive geometric image transformations between different views of the same scene or the same spatio-temporal event.
Irrespective of the validity of the stated biological predictions with regard to possible future more detailed neurophysiological evidence, the presented theory for covariant receptive fields under geometric image transformations reveals the structure of the close interaction between geometric image transformation and receptive field responses that holds also if a vision system would choose possible other ways of computing the receptive responses, for example according to a corresponding multi-parameter Lie group based on a set of infinitesimal generators over each degree of freedom for the set of primitive geometric image transformations.

### v4

@@ 6.2 Additional variabilities involving the temporal domain
Concerning the degree of freedom corresponding to temporal scaling transformations, the corresponding degree of freedom in terms of the temporal scale parameter $\sigma_{t}=\sqrt{\tau}$ is also special in the sense that both the non-causal temporal Gaussian kernel and time-causal limit kernel used for temporal smoothing in the idealized model (4.3) for spatio-temporal receptive fields obey cascade properties over temporal scales of the form (Lindeberg 2023a Equation (2))
$h(\cdot;\;\tau_{2})=(\Delta h)(\cdot;\;\tau_{1}\mapsto\tau_{2})*h(\cdot;\;\tau_{1})$ (132)
for any pair of temporal scales $(\tau_{1},\tau_{2})$ with $\tau_{2}>\tau_{1}$ and for some family of transformation kernels $(\Delta h)(t;\;\tau_{1}\mapsto\tau_{2})$. This implies that any receptive field responses, here represented as the temporal derivative $L_{t^{n}}(\cdot;\;\tau)$ at any coarser temporal scale $\tau_{2}$ can be computed by an additional temporal filtering operation being applied to the receptive field responses, here represented as the temporal derivatives $L_{t^{n}}(\cdot;\;\tau)$ at any finer temporal scale $\tau_{1}$, according to
$L_{t^{n}}(\cdot;\;\tau_{2})=(\Delta h)(\cdot;\;\tau_{1}\mapsto\tau_{2})*L_{t^{n}}(\cdot;\;\tau_{1}).$ (133)
In the case of a non-causal temporal domain, the difference kernel $\Delta h$ is a Gaussian kernel $(\Delta h)(t;\;\tau_{1}\mapsto\tau_{2})=g(t;\;\tau_{2}-\tau_{1})$ because of the semi-group property of the Gaussian kernel. In the case of a time-causal domain, the difference kernel is either a single truncated exponential kernel $h_{\exp}(t;\;\mu_{k})=1/\mu_{k}\exp(-t/\mu_{k})$ if $t>0$, or 0 otherwise, or the convolution of a set of such truncated exponential kernels, with appropriate time constants $\mu_{k}$ to fit the relationship between the temporal scale levels $\tau_{1}$ and $\tau_{2}$.
This, in turn, means that a vision system could, in principle, choose to only implement the earliest layers of temporal receptive fields at the finest temporal scale and nevertheless have the ability to compute the representations at coarser temporal scales, based on additional temporal smoothing applied to the temporal or spatio-temporal receptive field representations at the finest temporal scales. Thus, irrespective of whether the temporal receptive fields are expanded over the temporal scales, it seems plausible that the vision system should have the ability to compute visual operations corresponding to temporal scale covariance.
@@
An extended theory for how an idealized vision system, based on receptive fields in terms of the generalized Gaussian derivative theory for visual receptive fields, could compute receptive field responses at coarser spatial and temporal scales, based on cascade smoothing properties from finer spatial and temporal scales, is given in Lindeberg (2025g).
@@
Given the above theoretically motivated predictions, main questions for further research concern:
  • For which of the above variabilities can neurophysiological support be found regarding expansions of the shapes of the receptive fields over the corresponding degrees of freedom for different types of species?
  • Are the degrees of freedom corresponding to spatial scaling transformations and temporal scaling transformations special, because of the properties that the smoothing operations over the spatial and the temporal scale parameters can be performed in terms of cascade smoothing operations, implying that only representations of explicit receptive field responses at the finest spatial and/or temporal scales would be needed to compute the receptive field responses at coarser and/or temporal scales?
    Does efficiency call for the visual system to only implement the finest spatial and temporal scales in the sensorium, or does the vision system implement redundancies in this respect?
  • If there would be expansions over both the degrees of freedom of the affine group beyond uniform scaling transformations and rotations as well as an expansion over the degrees of freedom of the Galilean group, are those expansions then performed jointly or separately? In other words, are such expansions performed over the tensor product of the corresponding affine and Galilean parameters or separately for each subset of these degrees of freedom?
    For efficiency reasons, separate expansions over the affine and the Galilean degrees of freedom would need much less wetware to span the degrees of freedom of the corresponding geometric image transformations. A joint expansion over these degrees of freedom could on the other hand allow for the computation of more accurate estimates of the structure of dynamic scenes.
  • How well do the idealized models for spatial and spatio-temporal receptive fields explicitly listed in Section 5.5 model the biological receptive fields for simple cells in the primary visual cortex for different species?
  • How wide ranges in the parameter spaces of the receptive fields are spanned for the respective degrees of freedom for different species?
Answering these questions could provide valuable cues about how well different types of species have adapted their visual systems to handle the influence of different types of geometric image transformations on the image data.
Unfortunately, previous neurophysiological studies of visual receptive fields may not have been designed to answer the questions raised from this theoretical background. Answering the question about variabilities in the shapes of the visual receptive fields corresponding to the simple cells in the primary cortex would call for the accumulation of shape parameters of reconstructed receptive fields for a relatively large number of visual neurons, as well as using visual stimuli that would probe against the listed types of variabilities with regard to the 4 main types of geometric image transformations as leading to a total number of 7 degrees of freedom.
Currently, it seems very hard to get access to raw reconstructed receptive field profiles from previous recordings of the visual receptive fields in V1 of higher mammals. Publicly available receptive fields, as published in scientific publications, typically only show a rather small number of visual receptive fields. Concerning the hypothesized variabilities with regard to the degree of elongation of visual receptive fields, the neurophysiological evidence, that we have based our previous hypothesis on, are based on orientation selectivity properties of higher mammals (primates or cats) with a clear pinwheel structure. Recordings from mice, for which there is the largest number of public data, may on the other hand not be able to fully answer this question, because the vision system of mice is far less developed compared the visual systems of higher mammals (see e.g. Huberman and Niell (2011)), and does specifically lack pinwheel structures. In addition, recent results by Fu et al. (2024) of reconstructing most exciting inputs (MEIs) for early visual receptive fields in mice suggest that complex spatial features emerge earlier in the visual pathway of mice compared to primates, and that the receptive fields of mice are therefore more complex than a Gabor or Gaussian derivative model, as used in the theoretical model for stating the predictions in this paper.
For these overall reasons, in Lindeberg (2023b) Sections 3.2.1–3.2.2 and Lindeberg (2025c) Sections 4.2–4.3, sets of suggestions for more specific neurophysiological experiments are proposed to investigate the underlying hypotheses in more detail, and to characterize the structure of biological receptive fields in the primary visual cortex with respect to the influence of parameters of the receptive fields corresponding to the different degrees of freedom of the 4 main types of geometric image transformations. Complementary suggestions for further research in relation to the influence of geometric image transformations on early vision are also outlined in Lindeberg (2025a) Section 8.1.
@@ 7 Summary and discussion
This theory is based on idealized models of visual receptive fields in terms of combinations of smoothing with spatial smoothing kernels of the form (26) or spatio-temporal smoothing kernels of the form (25) with scale-normalized spatial and temporal derivatives according to Section 5.3. In Sections 5.4 and 5.5, we have described how the resulting idealized models of spatial or spatio-temporal receptive fields obey provable covariance properties under compositions of spatial scaling transformations, spatial affine transformations, Galilean transformations and temporal scaling transformations. Specifically, we have in Section 6 considered the hypothesis about whether the receptive fields of simple cells in the primary visual cortex can be regarded as having their shapes expanded with regard to the degrees of freedom of the basic types of geometric image transformations that occur in the image formation process.
By postulating that the responses of idealized models of receptive fields in terms of scale-normalized spatial and temporal derivative operators should be possible to match between the image domains before and after the geometric image transformations, we have predicted a set of variabilities over (i) spatial scaling transformations, (ii) image rotations, (iii) the degree of spatial elongation of the receptive fields, (iv) a 4:th spatial degree of freedom, (v) Galilean transformation over image space-time and (vi) temporal scaling transformations. We have considered potential support for the covariance properties of the either purely spatial or joint spatio-temporal receptive fields with regard to these 7 degrees of freedom (the Galilean transformation comprises 2 degrees of freedom), in view of neurophysiological evidence and structural properties regarding how populations of receptive field responses can be computed based on structural properties of the families of receptive fields under variabilities over the filter parameters. In the absence of sufficient neurophysiological or psychophysical evidence to firmly state whether the predicted properties would hold in the primary visual cortex of higher mammals, we have in Section 6.3 pointed to directions for future neurophysiological investigations to characterize these properties in more detail.
@@
  • the spatial scaling factor $S_{t}$ corresponds to the inverse depth $1/Z$, if the affine transformation matrix $A$ in the composed monocular image transformation (5) is normalized in such a way that the affine transformation matrix $A$ reflects a scaled orthographic projection model,
@@
Hence, a vision system that is able to extract these parameters of the geometric image transformations based on processing and comparing populations of receptive field responses, should in principle have the ability to compute direct cues to the structure of environment, directly from established matching relations over the receptive field responses between the different views, or in relation to a learned memory of receptive field responses from previous views. Thereby, certain functionalities of the vision system could be formulated directly in terms of the parameters of the primitive geometric image transformations between different views of the same scene or the same spatio-temporal event.
Thus, in summary, we have presented a coherent theory for how the structure of the geometric image transformations, that determine the formation of the image data that reach the visual sensors in the perception system, influence the desired structure of the receptive field responses in an idealized vision system. Specifically, we have argued how we based on the requirement of formulating identity relations between receptive field responses acquired under different viewing conditions for the same scene, are lead to the requirement of covariance properties of the receptive fields, which in turn imply a set of variabilities of the shapes of the receptive fields, to span the degrees of freedom of the corresponding geometric image transformations. In this way, we have formulated a set of theoretical predictions, and also discussed partial support of these predictions in view of existing neurophysiological evidence and complementary theoretical analysis in terms of cascade smoothing properties.
Finally, concerning the related sense of hearing, briefly touched upon in the beginning of the introduction, let us state that there is a way to define corresponding provably covariant receptive fields for auditory signals, based on the theories in Lindeberg and Friberg (2015) and Lindeberg (2025f), see specifically Section X in the latter reference.